Length of a rectangular field is 250m and width is 150m. Anuradha runs around this field 3 times. How far did she run? How many times she should run around the field to cover a distance of 4km?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

From the question it is given that,
Length of a rectangular field is 250m
Width of rectangular field = 150 m
Perimeter of rectangle = 2 × (length + breadth)

= 2 × (250 + 150) = 2 × 400
= 800 m

Given that, Anuradha runs around this field 3 times = 3 × 800 = 2400 m

We know that
1km = 1000 m
So
2400/1000
= 2.4 km i.e. 2 km 400 m
Number of times Anurudha should run around the field to cover a distance of 4km = 4000/800

= 5 times
Therefore, 5 times Anurudha should run around the field to cover a distance of 4km.


Thanks